OpenStudy (anonymous):

A quantity of energy equal to 8.90 x 10^12 joules was emitted from a nuclear reaction. What amount of mass was required to be converted to yield this amount of energy? (1 J = 1 kg m2/s2) Answers: 9.90 x 10^-5 kg 2.96 x 10^-4 kg 2.67 x 10^4 kg 8.01 x 10^5 kg

OpenStudy (chmvijay):

E=MC^2 M =E/C^2 where e is energy c is velcity of light in vacuum M= mass find it

OpenStudy (anonymous):

I tried that, but I got 1.008988764 x 10^28. just to be sure, e= 8.90 x 10^12, and c^2= 8.98 x 10^16, right?

OpenStudy (chmvijay):

M= 8.90 * 10^12 / (3 *10^8)^2 =8.90*10^12 /9*10^16 =0.98888*10^12-16 =0.988*10^-4 =9.888*10^-5 =9.9*10^-5

OpenStudy (chmvijay):

LOL u used formula wrong i think :) M=E/C^2 is formula

OpenStudy (anonymous):

Thank you very much for you help!

OpenStudy (chmvijay):

y welcome